OpenStudy (anonymous):

Prove that \[4\sin(x+{1\over6}\pi)\sin(x-{1\over6}\pi)=3-4\cos^2x\]

OpenStudy (dumbcow):

sin(a+b) = sina*cosb + sinb*cosa sin(a-b) = sina*cosb - sinb*cosa a = x b = pi/6 sin b = 1/2 cos b = sqrt3/2

OpenStudy (anonymous):

Yes, I know that. But how do I work with the 4sin?

OpenStudy (callisto):

4sin(x+1/6π)sin(x−1/6π) = -2{ cos[(x+1/6π)+(x-1/6π)] - cos[(x+1/6π)-(x-1/6π)] } = 2cos 1/3π -2 cos 2x = 1 - 2[ 2(cosx)^2 -1] = 1- 4(cosx)^2 +2 = 3-4(cosx)^2

OpenStudy (dumbcow):

\[\rightarrow 4(\frac{\sqrt{3}}{2}\sin x+\frac{1}{2} \cos x)(\frac{\sqrt{3}}{2}\sin x-\frac{1}{2} \cos x)\] \[=4(\frac{3}{4}\sin^{2} x - \frac{1}{4}\cos^{2} x)\] \[=3 \sin^{2} x - \cos^{2} x\]

OpenStudy (callisto):

first step : product to sum formula second step: double angle formula

OpenStudy (anonymous):

@Callisto How? I don't understand...

OpenStudy (callisto):

sinAsinB = -1/2 [cos(A+B)-cos(A-B) ] put A= x+1/6π and B = x-1/6π then you'll get 2cos 1/3π -2 cos 2x

OpenStudy (callisto):

then use double angle formula, cos2x = 2(cosx)^2 -1 to reduce 2cos 1/3π -2 cos 2x ps : 2cos 1/3π = 2 (0.5) =1 -2 cos 2x = -2[2(cosx)^2 -1] = -4(cosx)^2 +2 combine the two, which gives 1-4(cosx)^2 +2 that is 3-4(cosx)^2

OpenStudy (anonymous):

I've never learned the product formula...? Is there another way to do it without the product formula?

OpenStudy (callisto):

hmm dumbcow has shown you lol only left one step for you that is 3(sinx)^2 - (cosx)^2 = 3 [ 1-(cosx)^2 ] - (cosx)^2 = 3 - 4 (cosx)^2

OpenStudy (anonymous):

how does 3(sinx)^2 become 3[ 1- (cosx)^2]?

OpenStudy (anonymous):

\[\sin^2x+\cos^2x=1\] so \[\sin^2x=1-\cos^2x\]

OpenStudy (anonymous):

Oh Ok. Thanks.